(Complex annuity​) Upon graduating from college 35 years​ ago, Dr. Nick Riviera was already thinking of retirement. Since then he has made deposits into his retirement fund on a weekly basis in the amount of ​$25. Nick has just completed his final payment and is at last ready to retire. His retirement fund has earned 7.1​% interest compounded weekly. a. How much has Nick accumulated in his retirement​ account? b. In addition to all​ this, 20 years​ ago, Nick received an inheritance check for ​$30 comma 000 from his beloved uncle. He decided to deposit the entire amount into his retirement fund. What is his current balance in the​ fund?
Answer #1

1.
=25/(7.1%/52)*((1+7.1%/52)^(52*35)-1)=201056.832057028

2.
=25/(7.1%/52)*((1+7.1%/52)^(52*35)-1)+30000*(1+7.1%/52)^(52*15)=288018.808913552